OpenStudy (anonymous):

Please check my answer. I got option D. @ganeshie8

OpenStudy (anonymous):

@ganeshie8 Please check work/answer! :)) ---------------------------- Volume=lbh Each dimension of the second prism is 4 times the first prism meaning Volume=(4l)(4b)(4h)=64lbh so it'll have 64 times more water than the first tank ! ~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~

ganeshie8 (ganeshie8):

Excellent !!

OpenStudy (anonymous):

Do you think you can help me out with one more? :/

ganeshie8 (ganeshie8):

sure, ask

OpenStudy (anonymous):

It's this one:

OpenStudy (anonymous):

Thanks :)

ganeshie8 (ganeshie8):

similar to previous question : multiply the volume by cube of scale factor

ganeshie8 (ganeshie8):

scale factor = 10/5 = 2

ganeshie8 (ganeshie8):

volume of second box = 52*2^3 = ?

OpenStudy (anonymous):

416 !

OpenStudy (anonymous):

My final answer would be ** option D. 416 ft^3 ?

ganeshie8 (ganeshie8):

Correct !

ganeshie8 (ganeshie8):

surface area : multiply by \(k^2\) voolume : multiply by \(k^3\)

ganeshie8 (ganeshie8):

the first example which you have worked earlier tells us why the multiplying business works

OpenStudy (anonymous):

Thank you sooo much! :DD

ganeshie8 (ganeshie8):

you're welcome :)